# Koopman Invariants as Drivers of Emergent Time-Series Clustering in Joint-Embedding Predictive Architectures

### v1

@@ 1 Introduction
Self-Supervised Learning (SSL) has emerged as a powerful paradigm for learning rich data representations from unlabeled data, driving significant progress across diverse domains (Chen2020SimpleFS; Grill2020BootstrapYO; Chen2021). These methods learn by solving pretext tasks, with the hope that the learned representations capture underlying structures of the data. However, despite their empirical power, the precise mechanisms by which some SSL architectures discover these structures often remain opaque, treating critical model components as black boxes. This underscores a pressing need for theoretical frameworks that can provide a principled understanding of their behavior and guide future development.
Among the diverse SSL strategies, Joint-Embedding Predictive Architectures (JEPAs) (LeCun2022PathToAutonomousAI; Assran2023SelfSupervisedLA) offer a compelling approach. Rather than reconstructing raw inputs, JEPA is a non-generative approach that learns by predicting future data representations within an abstract, jointly-embedded latent space. This focus on abstract predictability has proven highly effective across diverse domains, spanning from static images to dynamic video and time-series data (verdenius2024latpfnjointembeddingpredictive; Assran2023SelfSupervisedLA; bardes2024revisiting).
@@
To address this, we turn to dynamical systems theory, specifically the Koopman operator framework (Koopman1931HamiltonianSA; Mezic2005SpectralPO). The Koopman operator offers a powerful way to analyze nonlinear dynamical systems by lifting observations into a space where their evolution becomes linear. This approach has inspired various machine learning techniques aiming to learn these linear representations. For example, significant research has focused on using deep autoencoders to learn intrinsic coordinates where dynamics evolve linearly under a learned Koopman operator, often with auxiliary networks to enforce desired properties like parsimony or to capture continuous spectra (Lusch2018; Yeung2019). Architectures like Linearly Recurrent Autoencoder Networks explicitly constrain latent dynamics to follow a linear recurrent layer, effectively learning finite-dimensional Koopman approximations (Azencot2020; Otto2019). This philosophy of enforcing linear latent dynamics has achieved remarkable success in modern structured State-Space Models (SSMs) like Mamba (gu2022efficientlymodelinglongsequences; gu2024mambalineartimesequencemodeling), which represent the current state-of-the-art for many sequence modeling tasks. These methods, building upon foundations like Dynamic Mode Decomposition and its extensions (DMD2010Schmid; Korda2018), demonstrate the utility of Koopman theory for system identification and representation learning where linear evolution is a primary target.
Other approaches, such as VAMPnets (VAMPNets2018Mardt) and Time-Lagged Autoencoders (Wehmeyer2018), are designed to learn slow collective variables or eigenfunctions of the underlying system’s transfer operator, which are crucial for understanding long-timescale dynamics and regime transitions. A distinct but related goal in representation learning is to achieve disentanglement, where latent variables are explicitly regularized to capture independent factors of variation in the data (higgins2017betavae; Kim2018). While these methods successfully identify key dynamical or generative modes, they often employ specialized objectives directly tied to these targets.
In contrast to these approaches that explicitly model system dynamics, optimize for specific eigenfunctions, or regularize for disentangled factors, the mechanism within JEPA is implicit. While many self-supervised methods for time series rely on contrastive objectives that learn similarity based on temporal proximity (Yue2022), we hypothesize that JEPA’s purely predictive objective in a latent space is what drives it to learn functions that are invariant under the system’s evolution within distinct dynamical regimes.
@@ 2 Theoretical Framework
The momentum coefficient $\alpha$ is typically chosen to be very close to 1 (e.g., 0.996 to 0.999), ensuring that $\theta_{EMA}$ represents a slowly evolving, stable average of $\theta$ (Grill2020BootstrapYO; He2020). The target encoder provides stable targets $z_{t+\Delta}=f_{\mathrm{EMA}}(x_{t+\Delta})$ for training.
@@
To perform a tractable theoretical analysis of the representations $f_{\theta}$ JEPA learns, we make a key simplifying assumption. We assume that the target encoder $f_{\mathrm{EMA}}$ closely tracks the online encoder $f_{\theta}$. Specifically, if the rate of change of the online parameters $\theta$ per training step is small relative to $(1-\alpha)$, then the difference $\|\theta-\theta_{EMA}\|$ remains small. Consequently, for well-behaved neural network functions $f$, this implies $f_{\mathrm{EMA}}(y)\approx f_{\theta}(y)$ for any relevant input $y$. This approximation is common in the analysis of similar self-supervised methods employing EMA targets (Grill2020BootstrapYO; He2020; tian2021understanding) and is implicitly validated by their empirical success, which relies on effective target stabilization.
@@ 4 Experimental Validation
(a) JEPA with MLP Predictor: Emergent, regime-aligned clustering.
(b) Autoencoder: Failure to separate regimes.
Figure 3: Latent space visualization (t-SNE) of test set embeddings. (a) JEPA embeddings often form distinct clusters corresponding to underlying dynamical regimes (indicated by colors). (b) Embeddings from a conventional autoencoder with an identical encoder architecture may not exhibit such clear regime-based separation on the same data.
@@ 5 Results and Discussion
A primary outcome of our theory is that JEPA’s encoder $f_{\theta}$ should learn representations $z_{t}=f_{\theta}(x_{t})$ that distinguish the underlying dynamical regimes, leading to distinct clusters in the latent space. To investigate this, we projected the $k$-dimensional latent embeddings of test set windows into a 2D space using t-SNE (vanderMaaten2008). The points were then color-coded according to their ground-truth regime labels.
@@ Appendix A Detailed Proofs for Section 3
(d) The functions $\{\chi_{1},\dots,\chi_{r}\}$ are linearly independent in $L^{2}(\mu)$ because their supports $\mathcal{X}_{i}$ are essentially disjoint and $\mu(\mathcal{X}_{i})=\alpha_{i}>0$. From part (a), they are all eigenfunctions of $\mathcal{K}$ with eigenvalue 1. Thus, the eigenspace corresponding to eigenvalue 1 has dimension at least $r$. A fundamental result in ergodic theory states that for a measure-preserving transformation, the dimension of the eigenspace of its Koopman operator corresponding to eigenvalue 1 is equal to the number of ergodic components of the invariant measure (Petersen1983ErgodicT; Eisner2015OperatorTE). Under Assumption 3.1, there are $r$ such ergodic components $(\mathcal{X}_{i},\mu_{i})$. Therefore, the space of functions in $L^{2}(\mu)$ that are invariant under $\mathcal{K}$ has dimension $r$. ∎
@@ Appendix B Synthetic Regime Generation Parameters
All ARMA processes are generated using statsmodels.tsa.arima_process.ArmaProcess (Seabold2010), ensuring stationarity for the chosen parameters.

### v2

@@ 1 Introduction
Self-Supervised Learning (SSL) has emerged as a powerful paradigm for learning rich data representations from unlabeled data, driving significant progress across diverse domains (Chen et al. 2020; Grill et al. 2020; Chen and He 2021). These methods learn by solving pretext tasks, with the hope that the learned representations capture underlying structures of the data. However, despite their empirical power, the precise mechanisms by which some SSL architectures discover these structures often remain opaque, treating critical model components as black boxes. This underscores a pressing need for theoretical frameworks that can provide a principled understanding of their behavior and guide future development.
Among the diverse SSL strategies, Joint-Embedding Predictive Architectures (JEPAs) (LeCun 2022; Assran et al. 2023) offer a compelling approach. Rather than reconstructing raw inputs, JEPA is a non-generative approach that learns by predicting future data representations within an abstract, jointly-embedded latent space. This focus on abstract predictability has proven highly effective across diverse domains, spanning from static images to dynamic video and time-series data (Verdenius et al. 2024; Assran et al. 2023; Bardes et al. 2024).
@@
To address this, we turn to dynamical systems theory, specifically the Koopman operator framework (Koopman 1931; Mezić 2005). The Koopman operator offers a powerful way to analyze nonlinear dynamical systems by lifting observations into a space where their evolution becomes linear. This approach has inspired various machine learning techniques aiming to learn these linear representations. For example, significant research has focused on using deep autoencoders to learn intrinsic coordinates where dynamics evolve linearly under a learned Koopman operator, often with auxiliary networks to enforce desired properties like parsimony or to capture continuous spectra (Lusch et al. 2018; Yeung et al. 2019). Architectures like Linearly Recurrent Autoencoder Networks explicitly constrain latent dynamics to follow a linear recurrent layer, effectively learning finite-dimensional Koopman approximations (Azencot et al. 2020; Otto and Rowley 2019). This philosophy of enforcing linear latent dynamics has achieved remarkable success in modern structured State-Space Models (SSMs) like Mamba (Gu et al. 2022; Gu and Dao 2024), which represent the current state-of-the-art for many sequence modeling tasks. These methods, building upon foundations like Dynamic Mode Decomposition and its extensions (Schmid 2010; Korda and Mezić 2018), demonstrate the utility of Koopman theory for system identification and representation learning where linear evolution is a primary target.
Other approaches, such as VAMPnets (Mardt et al. 2018) and Time-Lagged Autoencoders (Wehmeyer and Noé 2018), are designed to learn slow collective variables or eigenfunctions of the underlying system’s transfer operator, which are crucial for understanding long-timescale dynamics and regime transitions. A distinct but related goal in representation learning is to achieve disentanglement, where latent variables are explicitly regularized to capture independent factors of variation in the data (Higgins et al. 2017; Kim and Mnih 2018). While these methods successfully identify key dynamical or generative modes, they often employ specialized objectives directly tied to these targets.
In contrast to these approaches that explicitly model system dynamics, optimize for specific eigenfunctions, or regularize for disentangled factors, the mechanism within JEPA is implicit. While many self-supervised methods for time series rely on contrastive objectives that learn similarity based on temporal proximity (Yue et al. 2022), we hypothesize that JEPA’s purely predictive objective in a latent space is what drives it to learn functions that are invariant under the system’s evolution within distinct dynamical regimes.
@@ 2 Theoretical Framework
The momentum coefficient $\alpha$ is typically chosen to be very close to 1 (e.g., 0.996 to 0.999), ensuring that $\theta_{EMA}$ represents a slowly evolving, stable average of $\theta$ (Grill et al. 2020; He et al. 2020). The target encoder provides stable targets $z_{t+\Delta}=f_{\mathrm{EMA}}(x_{t+\Delta})$ for training.
@@
To perform a tractable theoretical analysis of the representations $f_{\theta}$ JEPA learns, we make a key simplifying assumption. We assume that the target encoder $f_{\mathrm{EMA}}$ closely tracks the online encoder $f_{\theta}$. Specifically, if the rate of change of the online parameters $\theta$ per training step is small relative to $(1-\alpha)$, then the difference $\|\theta-\theta_{EMA}\|$ remains small. Consequently, for well-behaved neural network functions $f$, this implies $f_{\mathrm{EMA}}(y)\approx f_{\theta}(y)$ for any relevant input $y$. This approximation is common in the analysis of similar self-supervised methods employing EMA targets (Grill et al. 2020; He et al. 2020; Tian et al. 2021) and is implicitly validated by their empirical success, which relies on effective target stabilization.
@@ 4 Experimental Validation
Figure 3: Latent space visualization (t-SNE) of test set embeddings. Left: JEPA embeddings often form distinct clusters corresponding to underlying dynamical regimes (indicated by colors). Right: Embeddings from a conventional autoencoder with an identical encoder architecture may not exhibit such clear regime-based separation on the same data.
@@ 5 Results and Discussion
A primary outcome of our theory is that JEPA’s encoder $f_{\theta}$ should learn representations $z_{t}=f_{\theta}(x_{t})$ that distinguish the underlying dynamical regimes, leading to distinct clusters in the latent space. To investigate this, we projected the $k$-dimensional latent embeddings of test set windows into a 2D space using t-SNE (van der Maaten and Hinton 2008). The points were then color-coded according to their ground-truth regime labels.
@@ Appendix A Detailed Proofs for Section 3
(d) The functions $\{\chi_{1},\dots,\chi_{r}\}$ are linearly independent in $L^{2}(\mu)$ because their supports $\mathcal{X}_{i}$ are essentially disjoint and $\mu(\mathcal{X}_{i})=\alpha_{i}>0$. From part (a), they are all eigenfunctions of $\mathcal{K}$ with eigenvalue 1. Thus, the eigenspace corresponding to eigenvalue 1 has dimension at least $r$. A fundamental result in ergodic theory states that for a measure-preserving transformation, the dimension of the eigenspace of its Koopman operator corresponding to eigenvalue 1 is equal to the number of ergodic components of the invariant measure (Petersen 1983; Eisner et al. 2015). Under Assumption 3.1, there are $r$ such ergodic components $(\mathcal{X}_{i},\mu_{i})$. Therefore, the space of functions in $L^{2}(\mu)$ that are invariant under $\mathcal{K}$ has dimension $r$. ∎
@@ Appendix B Synthetic Regime Generation Parameters
All ARMA processes are generated using statsmodels.tsa.arima_process.ArmaProcess (Seabold and Perktold 2010), ensuring stationarity for the chosen parameters.